OpenStudy (anonymous):

Please solve this (^=means exponent) (x^3^p+1 y^p+2)

OpenStudy (anonymous):

Please make your equation clear.

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

x raise to 2 p plus1 y raise to plus2...

OpenStudy (anonymous):

x raise to 2 p plus1 y raise p to plus2...

OpenStudy (anonymous):

Oh. it's just a term

OpenStudy (anonymous):

what do u think is the answer?

OpenStudy (anonymous):

You cannot solve this in any way.

OpenStudy (anonymous):

apply it on the rules of exponent

OpenStudy (anonymous):

Your expression has no meaning.

OpenStudy (anonymous):

write exactly character, one by one, as you have them on youjr sheet

OpenStudy (anonymous):

your* sheet

OpenStudy (anonymous):

ok w8 a sec

OpenStudy (anonymous):

x

OpenStudy (anonymous):

^3p+1

OpenStudy (anonymous):

y

OpenStudy (anonymous):

You made another post with the same problem.

OpenStudy (anonymous):

no no... all togather... in the same line

OpenStudy (anonymous):

(x^3p+1 y^p+2)

OpenStudy (anonymous):

Just multiply the exponents inside the brackets by 3. eg 2^3 = 2 * 2 * 2 and (2^2)^3 = 2^2 * 2^2 *2^2

OpenStudy (anonymous):

if that's the case, we'll have: x^(9p+3)y^(3p+6)